tyā (x,y)メ(0,0) (x,y)= (0,0). if if

(e) Find the gradient of f(x, y) at the point (zo,o) (0,0)

(c) Find the equation of the tangent plane to the graph of the function z -f(x, y) at the point (x,y,z) (1,0,0).

Please do the parts in the given order
tyā (x,y)メ(0,0) (x,y)= (0,0). if if
1 (d) Given the unit vector u-( find the directional derivative of f(x, y) at the 리지, ,- point (to,m) = (0,0), in the direction of the vector a.
(e) Find the gradient of f(x, y) at the point (zo,o) (0,0)
(c) Find the equation of the tangent plane to the graph of the function z -f(x, y) at the point (x,y,z) (1,0,0).
Answer #1


f, (0,0) = lim ,f( 0 +1,0)-f(0,0)=limfih,0)f( 0,0 )=1 0-0 =0 0.0)= lim-f( 0,0 + k)-f(0,0)=lim 0-0-0 So the directional deriv